# Hybrid Deep Learning Model to Estimate Cognitive Effort from fNIRS Signals in Educational Game Playing

### v2

@@ 3.1.3. Cognitive Effort Measure
In this study, we aimed to estimate the RNE and RNI in each session segment. We determined the overall average and standard deviation (SD) for each session (segment 1 & 2 combined, segment 3 & 4 combined).
To calculate RNE and RNI, we took the average of $\Delta HbO$ to measure oxygenated hemoglobin concentrations in the blood during the learning period in the overall PFC.The $\Delta HbO$ and performance score values were converted to Z-scores to reduce individual variability and to bring effort and performance onto the same scale which makes comparisons and RNE/RNI calculations more interpretable. In this study, we used the mean score as performance ($P_{z}$) and the inverse mean of $\Delta HbO$ as cognitive effort ($CE_{z}$) because we want better performance with minimum cognitive effort (Paas et al., 2003; Shewokis et al., 2015; Getchell and Shewokis, 2023; Koiler et al., 2022; Sharmin et al., 2024a). The participants can answer all questions correctly. To avoid division by zero when the standard deviation of the performance score became zero (i.e., in cases of uniform performance), we added a small epsilon constant ($\epsilon=0.001$) to the denominator of the z-score equations. This ensured numerical stability and preserved correct directional trends in RNE and RNI calculations (Equation 4).
@@
[…]
@@ 3.2. Data Acquisition
We collected fNIRS data using COBI, and the fNIRS recording rate was 10 samples per second. The data was saved in .fNIRS format. Each question was presented for 30 seconds, so each file contained 300 rows for each question corresponding to 16 optode ($\Delta HbO$) values. Sixteen participants answered 16 questions each, leading to 256 responses in the dataset ($16\times 16$). The dataset was imbalanced because 168 responses were correct (class 1) and 88 were incorrect (class 0). This imbalance occurred because participants generally performed well, which suggests that task difficulty may need future adjustment for broader classification analyses. The dataset has 1228800 data points before pre-processing ($participants(16)\times questions(16)\times fNIRSpoint(300)\times optodes(16)$). Since fNIRS captures $\Delta HbO$) changes over time, the data is temporal in nature.
@@ 3.3. Data Pre-processing
After data acquisition, we pre-process our data using fNIRSoft software (see Figure 4). We checked raw light intensities and individual optodes visually.
Next, we used a finite impulse response filter (20th order, Hamming window), a low pass filter, which removed the raw light intensity data at 0.1 Hz input from physiological signals, such as respiration and heartbeat. Then we applied the Modified-Beer Lambert rule, where fNIRS transforms light intensity measurements into changes in the concentrations of $\Delta HbO$ (Cakir et al., 2015; Çakır et al., 2016). According to the rule, a light’s attenuation depends on the medium’s characteristics it passes through (Cakir et al., 2015; Çakır et al., 2016). A portion of the near-infrared light used in fNIRS is absorbed by hemoglobin in the blood arteries in the brain. oxygenated hemoglobin ($\Delta HbO$) concentrations are connected to changes in brain activity. fNIRS can quantify how much light is absorbed by calculating changes in $\Delta HbO$ (Cakir et al., 2015; Çakır et al., 2016). Finally, the detrending filter was applied to data characterizing changes in concentration to remove drift in the data using linear parameters that convert the baseline slope to zero. We got oxygenated hemoglobin $\Delta HbO$ from the available biomarkers. We labeled the data using fNIRSoft software and saved them in CSV file format.
Figure 4. Pre-processing and artifact removal of fNIRS data.
[…]
Figure 4: This figure shows the pre-processing steps to remove noise and artifact removal. In the first box, we checked raw light intensities and individual optodes visually. In the next text box, we used a finite impulse response filter (20th order, Hamming window), a low pass filter, where the raw light intensity data at 0.1 Hz input from physiological signals, such as respiration and heartbeat. For the third box, we applied the Modified-Beer Lambert rule, where fNIRS transforms measurements of light intensity into changes in the concentrations of $\Delta HbO$ and $\Delta HbO$. In the fourth box, the detrending filter was applied to data characterizing changes in concentration to remove drift in the data using linear parameters that convert the slope of the baseline to zero. In the fifth box, we labeled the fNIRS data based on task and time. and in the last box, we extract the $\Delta HbO$ in a .CSV file
Algorithm  LABEL:algo:_pre-process shows the steps of pre-processing used to prepare the fNIRS dataset for the proposed DeepNet model. The pipeline includes data extraction, missing value handling, standardization, feature selection, class balancing, reshaping, and train-test splitting. These steps ensure the dataset is clean, well-structured, and optimized for GRU-based model training.
At first, we extracted the quiz responses based on filenames. Then, we handled missing values by imputing with the mean per question and cleaned column names to prevent errors. Since participants often answered within 20 seconds, only the first 200 rows per question were retained. We grouped the data for performance (participant and question) prediction. After applying standardization (Z-score normalization) to ensure uniform feature scaling, we used Principal Component Analysis (PCA) to reduce dimensionality from 16 to 12 features (see LABEL:fig:PCA) to improve efficiency.
Finally, the dataset was split into 80% training and 20% testing. We had 256 quiz responses in total. After splitting, we had 208 responses (13 participants) for the training set and 48 (3 participants) for the test set. We had 134 data for correct answers (class 1) and 74 for wrong answers (class 0) in the training set. To balance this dataset, we applied the Synthetic Minority Over-Sampling Technique (SMOTE), which balances class distribution by oversampling the minority class. The data was then reshaped into (samples, 1 timestamp, 12 features) to be compatible with the GRU model.
@@ 4. Model Architecture
Our proposed hybrid DeepNet model (CNN-GRU) takes a reshaped PCA-processed input of shape (1, 12) per sample. The pipeline includes: A Conv1D layer with 32 filters and a kernel size of 1 to extract spatial patterns, a GRU layer with tunable hidden units (8 or 16) to capture temporal dependencies, a BatchNormalization and Dropout to reduce overfitting, and a fully connected dense layer of 64 units followed by a softmax output layer for binary classification. The CNN identifies which brain regions are most active during different cognitive states represented by fNIRS optodes. This helps the model to detect patterns related to effort.The GRU layer handles temporal dependencies in hemodynamic signals over the 30-second question window. This makes it well-suited to capture cognitive effort changes over time. This hybrid architecture is thus ideal for fNIRS data, which is both spatially distributed and temporally dynamic.
@@
Post-Training Analysis: Following model training, GRU embeddeddings were used as input features to an XGBoost classifier to further evaluate learned representations. For each model variant, we report Train Accuracy, Test Accuracy, Precision, Recall, F1-Score, and XGBoost Accuracy on the extracted features.
@@
GRU units: 8, 16; Dropout rates: 0.1, 0.2, 0.4; Learning rates: 0.0005, 0.001, 0.003; Batch sizes: 1, 4, 8, 16, 32.
@@
Figure 5. DeepNet Model Architecture. We fed fNIRS data into our model and to get performance score as output
[…]
Figure 6: DeepNet Model Architecture. The flow shows the dataset goes through the CNN layer, batch normalization, GRU layer, dropout(0.1), then dense layer, and finally the Softmax layer to get the final output.
@@
Table 1. Hyperparameters Used in the Model
[…]
Table 1: This table shows the Hyperparameter Used in the Model. It has two column parameters and its value. Hyperparameter Value Learning Rate 0.003 GRU Units 8 Convolutional Filters 32 Dropout Rate 0.1 Batch Size 4 Optimizer Adam Loss Function Categorical Cross-Entropy
@@ 5.1. Prediction of Performance Score
At first, we ran our dataset through traditional ML algorithms. Table 2 presents the evaluation metrics for different ML models based on 5-fold cross-validation. In most cases, we obtained higher training accuracy than test accuracy. The Random Forest model achieved the highest recall ($0.92$) and F1-score ($0.77$). SVM and XGBoost also demonstrated strong recall values ($0.82$ and $0.80$, respectively), with SVM achieving an F1-score of $0.73$. Gradient Boosting and CatBoost showed balanced precision-recall values, which indicates stable performance across categories. However, Neural Network and LightGBM showed poor generalization with the lowest test accuracy values ($0.65$ and $0.54$, respectively). Overall, models like Random Forest, SVM, and XGBoost exhibited higher sensitivity and robustness, while Neural Network and LightGBM struggled with consistency in classification. Also, the models were overfitted.
After evaluating traditional methods, we applied the dataset to various deep-learning models. Table 2 (bottom section) shows the performance comparison of different DeepNet architectures for the binary classification task. Our proposed hybrid model, CNN-GRU, outperformed all baseline models. It achieved a test accuracy of 73.08%, with a remarkable F1 score of 82.05%, demonstrating excellent performance in balancing precision (73.81%) and recall (91.18%). The notable improvement in recall indicates that the CNN-GRU model is particularly effective at minimizing false negatives.
@@
Table 2. Performance Comparison of Machine Learning and DeepNet Models
[…]
Table 2: compares the performance of machine and DeepNet models. the result shows that the proposed CNN-GRU model gives better accuracy to predict performance score. Model Train Test Precision Recall F1 Score ML Models Random Forest 1.00 0.63 0.66 0.92 0.77 Gradient Boosting 1.00 0.59 0.60 0.74 0.70 AdaBoost 0.98 0.57 0.66 0.73 0.69 SVM 0.70 0.64 0.65 0.82 0.73 Neural Network 0.71 0.65 0.63 0.70 0.66 LightGBM 1.00 0.54 0.59 0.73 0.68 CatBoost 1.00 0.64 0.65 0.81 0.72 DL Models CNN 0.68 0.56 0.56 0.56 0.56 LSTM 0.68 0.60 0.61 0.60 0.60 BiLSTM 0.72 0.66 0.66 0.66 0.66 CNN-GRU* 0.78 0.73 0.74 0.91 0.82
We conducted paired t-tests comparing CNN-GRU with each model using 5-fold test accuracies. While CNN-GRU performed better than others, none of the differences were statistically significant ($p>0.05$). Yet, these findings indicate that DeepNet models like the proposed CNN-GRU hybrid offer better generalization and classification capability for this time series task. This also suggests that learning directly from data using deep models provides better predictive performance compared to traditional approaches.
@@ 5.2. Feature Attribution Analysis (Interpretability)
To investigate which brain regions most influenced the model’s predictions, we conducted a multi-stage interpretability analysis combining PCA, GRU feature correlation, and SHAP explanations. 1. PCA-Optode and Region Mapping: we aggregated optode contributions by anatomical region LPFC and VMPFC to obtain regional level insight. PC2 exhibited a stronger total contribution from LPFC $(6.99)$ compared to VMPFC $(5.36)$, which suggests that this component reflects distributed LPFC activity.
2. GRU - PCA Correlation (Feature Encoding Stage) We then analyzed how GRU latent features (from the encoding stage) relate to the PCA components. As shown in Figure 6, GRU Feature 6 had the highest correlation with PC2 $(r=0.65)$. Given PC2 was previously found to be LPFC-dominated, this suggests GRU6 is encoding task-relevant signals originating from LPFC.
Figure 6. Correlation between GRU features and PCA components
3. SHAP - GRU Feature Importance (Decision Stage) Finally, we applied SHAP (SHapley Additive exPlanations) to identify which GRU features most influenced the model’s final predictions. As shown in Figure 7, GRU Features 3 and 0 had the highest SHAP values, indicating their strong impact in decision-making. Interestingly, these features also showed notable correlation with PC2—GRU3 negatively $(r=-0.67)$ and GRU0 positively $(r=0.42)$.
Figure 7. SHAP summary plot showing GRU feature importance in prediction
@@ 5.3. Estimation of Cognitive Effort
After identifying the proposed CNN-GRU as the best-performing model, we used its predicted performance scores to calculate RNE and RNI. These metrics provided a deeper understanding of participants’ cognitive states across task segments. Our test set included three participants: P8, P11, and P16.
Table 3 presents a detailed comparison of actual and predicted RNE and RNI values across task segments for three test participants. We calculated the RNE and RNI from the predicted score from the deep learning model ( Equation 4). The final row shows statistical similarity using Mean Absolute Error (MAE) and Pearson correlation. The very low MAE and near-perfect correlation $(r=0.999)$ confirm that the model’s predicted RNE and RNI values closely follow the actual measurements.
(a) RNE Comparison
(b) RNI Comparison
Figure 8. Line plots of predicted and actual RNE and RNI values across task segments for each test case.
[…]
Figure 7: Line plots of predicted and actual RNE and RNI values across task segments for each test case. The lines show that the predicted line follows the actual one.
Table 3. Comparison of Actual and Predicted RNE and RNI across Participants and Task Segments. Predicted values closely follow the actual values, as confirmed by low MAE and high Pearson correlation.
[…]
Table shows actual and predicted values of RNE and RNI for participants P8, P11, and P16 across 4 segments. A summary row shows statistical similarity using Mean Absolute Error (MAE) and Pearson correlation coefficient. Participant Segment RNE RNI Actual Predicted Actual Predicted P8 1 0.14 0.9 -0.63 0.13 2 0.25 -0.51 0.25 -0.51 3 -32.65 -32.98 32.65 32.32 4 -0.0002 0.33 0.0002 0.33 P11 1 0.95 0.53 -0.61 -1.03 2 -0.17 0.25 -0.17 0.25 3 6.80 6.94 -7.57 -7.44 4 0.38 0.25 0.38 0.24 P16 1 -2.33 -2.46 -1.83 -1.70 2 0.25 0.38 -0.25 0.38 3 -0.59 -0.59 -0.59 -0.59 4 0.0002 0.0002 0.0002 0.0002 Mean absolute error (MAE) 0.29 0.38 Pearson Correlation 0.99 0.99
To understand participants efficiency and involvement we plot $CE$ in X axis and $P_{z}$ in Y axis (see  Equation 4). Figure LABEL:fig:RNE_RNI_ML_dynamic visualizes participant-wise relative neural efficiency (RNE) and involvement (RNI) in Cartesian quadrants representing four cognitive states: HE+HI (High Effort, High Involvement), HE+LI (High Effort, Low Involvement), LE+HI (Low Effort, High Involvement), and LE+LI (Low Effort, Low Involvement). Predicted coordinates from the DeepNet model were plotted alongside the actual values for three participants (P8, P11, P16) across four task segments.
Out of the 12 plotted points, we can see that 09 predicted coordinates matched the actual cognitive states. This highlights the accuracy and interpretability of our CNN-GRU-based performance prediction model for estimating RNE and RNI.
In segment 1, participants began in LE+HI or LE+LI regions, which suggests that their initial performance was high with low cognitive effort, which is considered a potentially favorable cognitive state for learning. In segment 2, while performance levels varied across participants, most points clustered around the origin, which indicates moderate cognitive effort and performance, and suggests a transitional cognitive state.
In segment 3, we noticed a drop in cognitive effort for P11 and P16. This could be the reason for the longer break after segment 2. P8 reported having a headache after segment 2, which could be the reason for increased effort in segment 3. However, in segment 4, the effort dropped after a short rest for P8. In the other two cases, in segment 4, participants maintained a general transition of cognitive state.

### v3

@@ 3.1.1. Cognitive Effort Measure
We designed the study with two sessions, each consisting of two segments. Between the two segments in each session, participants were given a 20-second rest period, while an 8–10 minute break separated the two sessions. Each segment contained 4 quiz questions, resulting in a total of 16 questions across the entire experiment.
Our objective was to estimate the participants’ cognitive effort during each segment. To achieve this, we used their fNIRS signals to predict quiz scores by using various machine and deep learning models. Based on these predicted scores and corresponding brain activation levels, we computed the participants’ cognitive effort.
To calculate RNE and RNI, we took the average of $\Delta HbO$ to measure oxygenated hemoglobin concentrations in the blood during the learning period in the overall PFC. The $\Delta HbO$ and performance score values were converted to Z-scores to reduce individual variability and to bring effort and performance onto the same scale which makes comparisons and RNE/RNI calculations more interpretable. We used the mean score as performance ($P_{z}$) and the inverse mean of $\Delta HbO$ as cognitive effort ($CE_{z}$) because we want better performance with minimum cognitive effort (Paas et al., 2003; Shewokis et al., 2015; Getchell and Shewokis, 2023; Koiler et al., 2022; Sharmin et al., 2025). For each participant, we calculated the overall average (GM) and standard deviation (SD) of fNIRS signals for each session. The participants can answer all questions correctly. To avoid division by zero when the standard deviation of the performance score became zero (i.e., in cases of uniform performance), we added a small epsilon constant ($\epsilon=0.001$) to the denominator of the z-score equations. This ensured numerical stability and preserved correct directional trends in RNE and RNI calculations (Equation 4).
@@
(2) $\hskip 56.9055ptCE_{z}=\dfrac{\dfrac{1}{\Delta HbO_{i}}-\dfrac{1}{\Delta HbO_{\text{(GM)}}}}{\dfrac{1}{\Delta HbO_{\text{(SD)}}}}\hskip 28.45274pt$
@@ 3.2. Data Acquisition
We collected fNIRS data using COBI, and the fNIRS recording rate was 10 samples per second. Each question’s length was 30 seconds, so each file contained 300 rows of 16 optode ($\Delta HbO$) values. Since participants often answered within 20 seconds, only the first 200 rows per question were retained. We had total 256 responses in the dataset ($16\times 16$). The dataset was imbalanced because 168 responses were correct (class 1) and 88 were incorrect (class 0). This imbalance occurred because participants generally performed well, which suggests that task difficulty may need future adjustment for broader classification analyses. The dataset has ($participants(16)\times questions(16)\times fNIRSpoint(200)\times optodes(16)$) data points.
@@ 3.3. Data Pre-processing
After data acquisition, we pre-process our signal using fNIRSoft software. Using this software, we removed noisy channels and applied finite impulse response filter,a low pass filter, and detrending filter to data characterizing changes in concentration to remove drift in the data. At the end, we labeled the $\Delta HbO$ data using fNIRSoft software based on the biomarkers sent via Python code (Sharmin et al., 2024a; Sharmin and Abrar, 2025).
After pre-processed the signal, we extracted the quiz responses based on filenames. Then, we handled missing values by imputing with the mean per question and cleaned column names to prevent errors. We grouped the data for performance (participant and question) prediction. After applying standardization (Z-score normalization) to ensure uniform feature scaling, we used Principal Component Analysis (PCA) to reduce dimensionality from 16 to 12 features to improve efficiency.
Finally, instead of applying k-fold, we split the dataset participant wise because the dataset is small (Khosla et al., 2019). We had 13 participants with 208 responses for the training set and 3 participants with 48 responses for the test set. We had 134 data for correct answers (class 1) and 74 for wrong answers (class 0) in the training set. To balance this dataset, we applied the Synthetic Minority Over-Sampling Technique (SMOTE), which balances class distribution by oversampling the minority class. The data was then reshaped into (samples, 1 timestamp, 12 features) to be compatible with the GRU model. SMOTE was only applied to the training set, while the test set was kept in its original, imbalanced form to ensure the evaluation reflects real-world data distribution.
@@ 4. Model Architecture
Our proposed hybrid DeepNet model (CNN-GRU) takes a reshaped PCA-processed input of shape (1, 12) per sample. The pipeline includes: A Conv1D layer with 32 filters and a kernel size of 1 to extract spatial patterns, a GRU layer with tunable hidden units (8 or 16) to capture temporal dependencies, a BatchNormalization and Dropout to reduce overfitting, and a fully connected dense layer of 64 units followed by a softmax output layer for binary classification. The CNN identifies which brain regions are most active during different cognitive states represented by fNIRS optodes. This helps the model to detect patterns related to effort. The GRU layer handles temporal dependencies in hemodynamic signals over the 20-second question window. This makes it well-suited to capture cognitive effort changes over time. This hybrid architecture is thus ideal for fNIRS data, which is both spatially distributed and temporally dynamic.
@@
Post-Training Analysis: Following model training, GRU was used as input features to an XGBoost classifier to further evaluate learned representations. For each model variant, we report Train Accuracy, Test Accuracy, Precision, Recall, F1-Score, and XGBoost Accuracy on the extracted features.
@@
GRU units: 8, 16;
Dropout rates: 0.1, 0.2, 0.4;
Learning rates: 0.0005, 0.001, 0.003;
Batch sizes: 1, 4, 8, 16, 32.
@@
Figure 2. DeepNet Model Architecture. We fed fNIRS data into our model and to get performance score as output:
@@
Table 1. Hyperparameters Used in the Model: :
Hyperparameter Value
Learning Rate 0.003
GRU Units 8
Convolutional Filters 32
Dropout Rate 0.1
Batch Size 4
Optimizer Adam
Loss Function Categorical Cross-Entropy
@@ 5.1. Prediction of Performance Score
At first, we ran our dataset through traditional ML algorithms. Table 2 presents the evaluation metrics for different ML models based on 5-fold cross-validation. The Random Forest model achieved the highest recall ($0.92$) and F1-score ($0.77$). SVM and XGBoost also demonstrated strong recall values ($0.82$ and $0.80$, respectively), with SVM achieving an F1-score of $0.73$. Gradient Boosting and CatBoost showed balanced precision-recall values, which indicates stable performance across categories. However, Neural Network and LightGBM showed poor generalization with the lowest test accuracy values ($0.65$ and $0.54$, respectively). Overall, models like Random Forest, SVM, and XGBoost exhibited higher sensitivity and robustness, while Neural Network and LightGBM struggled with consistency in classification.
After evaluating traditional methods, we applied the dataset to various deep-learning models. Table 2 (bottom section) shows the performance comparison of different DeepNet architectures for the binary classification task. Our proposed hybrid model, CNN-GRU, outperformed all baseline models. It achieved an accuracy of 73.08%, with a remarkable F1 score of 82.05%, demonstrating excellent performance in balancing precision (74%) and recall (91.18%). The notable improvement in recall indicates that the CNN-GRU model is particularly effective at minimizing false negatives.
@@
Table 2. Performance Comparison of Machine Learning and DeepNet Models: :
Model Accuracy Precision Recall F1 Score
ML Models
Random Forest 0.63 0.66 0.92 0.77
XGBoost 0.59 0.60 0.74 0.70
AdaBoost 0.57 0.66 0.73 0.69
SVM 0.64 0.65 0.82 0.73
Neural Network 0.65 0.63 0.70 0.66
LightGBM 0.54 0.59 0.73 0.68
CatBoost 0.64 0.65 0.81 0.72
DL Models
CNN 0.56 0.56 0.56 0.56
LSTM 0.60 0.61 0.60 0.60
BiLSTM 0.66 0.66 0.66 0.66
CNN-GRU* 0.73 0.74 0.91 0.82
@@ 5.2. Feature Attribution Analysis (Interpretability)
To investigate which brain regions most influenced the model’s predictions, we conducted a multi-stage interpretability analysis combining PCA, GRU feature correlation, and SHAP explanations. 1. PCA-Optode and Region Mapping: We aggregated optode contributions by anatomical region Lateral PFC (LPFC) (optode 1-4; optode 13-16) and Ventromedial PFC (VMPFC) (optode 5-12) to obtain regional level insight. The LPFC is linked to goal representation, working memory, visuospatial attention, adaptive behavior, and cognitive control. In contrast, the VMPFC is involved in value-based decision-making, reward anticipation, and self-related evaluation (Sharmin et al., 2024b; Hasan et al., 2023; Getchell and Shewokis, 2023). PC2 exhibited a stronger total contribution from LPFC $(6.99)$ compared to VMPFC $(5.36)$, which suggests that this component reflects distributed LPFC activity.
2. GRU - PCA Correlation (Feature Encoding Stage): We then analyzed how GRU latent features (from the encoding stage) relate to the PCA components. As shown in Figure 3, GRU Feature 6 had the highest correlation with PC2 $(r=0.65)$. Given PC2 was previously found to be LPFC-dominated, this suggests GRU6 is encoding task-relevant signals originating from LPFC.
Figure 3. Correlation between GRU features and PCA components:
3. SHAP - GRU Feature Importance (Decision Stage) Finally, we applied SHAP (SHapley Additive exPlanations) to identify which GRU features most influenced the model’s final predictions. As shown in Figure 4, GRU Features 3 and 0 had the highest SHAP values, indicating their strong impact in decision-making. Interestingly, these features also showed notable correlation with PC2—GRU3 negatively (r=0.67)(r=-0.67) and GRU0 positively $(r=0.42)$.
Figure 4. SHAP summary plot showing GRU feature importance in prediction:
@@ 5.3. Estimation of Cognitive Effort
To evaluate the reliability of predicted cognitive effort metrics, we computed RNE and RNI from the predicted scores and compared them with actual RNE/RNI values across 12 task segments for three representative participants (P8, P11, and P16).
As shown in  Table 3, participant-wise Mean Absolute Error (MAE) and Pearson correlation were calculated between actual and predicted values. The predicted RNE and RNI metrics exhibited low error (MAE range: $0.065$ to $0.545$) and high correlation with actual values (Pearson $r>0.94$ for all cases), confirming that predicted performance can reliably approximate cognitive effort trends.  Figure 5 shows scatter plots comparing actual vs. predicted values of RNE (left) and RNI (right) across all segments (12 points each). The diagonal y = x line represents perfect prediction. The majority of points closely cluster around this line, indicating that the predicted metrics are highly accurate. This visual alignment is supported quantitatively by a high Pearson correlation (r = 0.99) and low Mean Absolute Error (RNE MAE = 0.29; RNI MAE = 0.38).
Table 3. Participant-wise MAE and Pearson Correlation for Predicted RNE and RNI:
# MAE Pearson
Participants RNE RNI RNE RNI
P8 0.55 0.54 0.99 0.99
P11 0.27 0.28 0.99 0.99
P16 0.06 0.19 0.99 0.94
Figure 5. Scatter plots comparing actual vs. predicted values of RNE (left) and RNI (right). The dotted diagonal line indicates perfect prediction. Data points cluster tightly around the line, demonstrating strong predictive accuracy.: